OpenStudy (anonymous):

The population of a local species of flies can be found using an infinite geometric series where a1 = 940 and the common ratio is . Write the sum in sigma notation and calculate the sum (if possible) that will be the upper limit of this population.

OpenStudy (anonymous):

PLEASE HELP :(

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

What's the common ratio?

OpenStudy (kc_kennylau):

\[\Sigma=\frac{a_1}{1-r}\]

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

one fifth is the common ratio

OpenStudy (kc_kennylau):

Use the formula that I just gave you

OpenStudy (anonymous):

alright give me a sec

OpenStudy (anonymous):

whats r?

OpenStudy (kc_kennylau):

The common ratio

OpenStudy (kc_kennylau):

Okay I shouldn't have used \(\Sigma\) as the sum

OpenStudy (anonymous):

im completely lost im sorry

OpenStudy (kc_kennylau):

The sum \(S\) of a geometric series with the first term being \(a_1\) and the common ratio being \(r\) is given by the formula \(S=\dfrac{a_1}{1-r}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4699/5?

OpenStudy (kc_kennylau):

And if you have to write it in sigma notation, \(\displaystyle\Large\sum_{n=1}^\infty a_1r^{n-1}\)

OpenStudy (anonymous):

939.8

OpenStudy (kc_kennylau):

Nope, 939.8 isn't what I got, mind showing us your steps?

OpenStudy (anonymous):

um yea i divided 940 by 1 minus 1/5

OpenStudy (kc_kennylau):

It's 940 divided by (1 minus 1/5) not (940 divided by 1) minus 1/5 ...

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

1175?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

so wat do i do now?

OpenStudy (kc_kennylau):

Well you have done the second part

OpenStudy (kc_kennylau):

The first part would be writing it as \(\displaystyle\Large\sum_{n=1}^\infty 940\times\left(\frac15\right)^{n-1}\)

OpenStudy (anonymous):

ok i did that

OpenStudy (kc_kennylau):

Then you'd have done everything

OpenStudy (anonymous):

the sum is 1,175

OpenStudy (anonymous):

yea its the other one the one that u did but the sum is 1,175 thank u so much

OpenStudy (kc_kennylau):

the lower limit would be i=0

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

is there any way u could help me with my last one?

OpenStudy (kc_kennylau):

If that's another question please open a new post

OpenStudy (anonymous):

ok